OpenStudy (anonymous):

HELP PLEASE!! Given the equation y − 3 =1/2 (x + 6) in point-slope form, identify the equation of the same line in slope-intercept form. x − 2y = −12 y =1/2 x y = 1/2x + 6 y = 1/2x + 9

OpenStudy (anonymous):

Sorry you distribute the 1/2, y-3=1/2x-3 Add the 3 to both sides

OpenStudy (anonymous):

You answer is y=1/2x :)

OpenStudy (anonymous):

THANK YOU SO MUCH!! i get now! :D thank you :)

OpenStudy (anonymous):

Your welcome can I get a medal :D

OpenStudy (anonymous):

Yay Thank U :)

OpenStudy (anonymous):

Welcome! :D its the least i could do thank you so much

OpenStudy (anonymous):

Ur Welcome :)